OpenStudy (thatonegirl_):

When is the slope of this = to 0?

OpenStudy (thatonegirl_):

\[\frac{ x^2+2x-8 }{ x+6 }\]

OpenStudy (thatonegirl_):

\[f'(x)=\frac{ 2x+2(x+6)-(x^2+2x-8) }{ (x+6)^2 }\]

OpenStudy (thatonegirl_):

\[2x+2x+12-x^2-2x+8=0\]

OpenStudy (thatonegirl_):

so then i got -(x^2-2x-20)=0 but how can i factor this?

OpenStudy (jackthegreatest):

ok you did f' right

OpenStudy (thatonegirl_):

sweet

OpenStudy (thatonegirl_):

where did i mess up?

OpenStudy (jackthegreatest):

+ 14x, not + 2x

OpenStudy (jackthegreatest):

you multiplied wrong

OpenStudy (jackthegreatest):

2x+14x+12−x2−2x+8=0

OpenStudy (thatonegirl_):

OMG I'm stupid af. I should've put that into parenthesis. Thank you! xD

OpenStudy (jackthegreatest):

np